OpenStudy (anonymous):

Identify the x-coordinate of any relative extrema of y = xln(x) , I know the answer need an explanation on approach thanks

OpenStudy (anonymous):

is it 1/e??

OpenStudy (anonymous):

Relative min at x = 1/e

OpenStudy (anonymous):

yeah, how can u get the maxima here??

OpenStudy (asnaseer):

you need to differentiate the function and then work out what value of x will it zero.

OpenStudy (asnaseer):

*will make it

OpenStudy (asnaseer):

@frank545 - do you know how to differentiate this function?

OpenStudy (slaaibak):

derivative using chain rule: f'(x) = x * 1/x + lnx = 1 + lnx 1 + lnx = 0 ln x = -1 x=e^-1=1/e

OpenStudy (anonymous):

Yes I understand differentiation, I was looking for a theoretical explanation in terms of what i look for when finding relative extrema

OpenStudy (asnaseer):

the differential define the slope of the tangent line to the curve

OpenStudy (asnaseer):

at a relative extrema, the slope is horizontal, and therefore equal to zero

OpenStudy (asnaseer):

e.g. look at this sample graph: